OpenStudy (anonymous):

If sin theta equals 4/5 and theta is in quadrant 2, the value of cot theta is?

OpenStudy (anonymous):

cos theta will be negative as it is in 2nd quadrant. cos^2 theta = 1- sin^2 theta = 1 - (4/5)^2 = 9/25 cos theta = - 3/5 cot theta = cos theta /sin theta = - 3/4

OpenStudy (anonymous):

THANK YOU SO MUCH!